Circular No. 8226

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003iu, 2003iv, AND 2003iw The discoveries of three apparent supernovae on unfiltered CCD survey

images have been reported — 2003iu by M. Armstrong (cf. IAUC 8207), 2003iv by J. Burket and W. Li (LOTOSS/KAIT; cf. IAUC 8224), and 2003iw by T. Boles (cf. IAUC 8212) and by Armstrong (tabulated data for SN 2003iw below from Boles).

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003iu Oct. 16.07 1^{h}55^{m}24.^{s}55 +21^{o}17′06′′.4 16.4 2′′ W, 3′′.6 S
2003iv Oct. 17.4 2 50 07.25 +12 50 46.5 17.2 23′′.2 W, 8′′.3 N
2003iw Oct. 17.95 21 39 43.03 + 6 17 03.0 17.6 22′′.0 W, 7′′.5 S

Additional approximate magnitudes by the respective discoverers (A = Armstrong, B = Boles, K = KAIT): SN 2003iu in UGC 1393, Sept. 1, [19.0 (A); 23, [19.5 (A); Oct. 17.106 UT, 16.4 (A). SN 2003iv in PGC 10738, Oct. 8.4, [19.0 (K); 18.4, 17.1 (K). SN 2003iw in NGC 7102, Sept. 15, [19.5 (B); 23, [19.5 (B); Oct. 7, [19.5 (B); 18.304, 17.6 (B); 18.873, 17.1 (A). Nothing is visible at the locations of SNe 2003iu or 2003iw on various Palomar Sky Survey plates. Armstrong adds that nothing was present at the location of SN 2003iw on his images from July 12 and Oct. 14, and he provides the following position end figures for the same object: 43^{s}.20, 02′′.3.

NOVAE IN M31 Further to IAUC 8222, M. Fiaschi, D. Tiveron, and A. Cardullo re-

port the discovery of an apparent nova located at α = 0^{h}42^{m}41^{s}.11, δ = +41^{o}16′16′′.1 (equinox 2000.0), which is 36′′.2 west and 7′′.6 north of the center of M31. Available Hα magnitudes from their images taken at the G. Colombo Observatory: Oct. 9 UT, [18; 15.865, 16.2; 16.998, 15.9.

Unfiltered CCD magnitudes by G. Sostero, Remanzacco, Italy, of Hornoch’s recent nova (IAUC 8222): Aug. 17, [18.5:; Sept. 29.83 UT, 16.7.

COMET C/2003 R1 (LINEAR)

Improved elliptical orbital elements from MPC 49768:

T = 2003 June 29.5876 TT  ω = 302.^{o}6817
e = 0.892795  Ω = 356.6908 }2000.0
q = 2.101742 AU  i = 149.1936

a = 19.604969 AU n^{o} = 0.0113542 P = 86.8 years

2003 October 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green